Find the slop and y-intercept of the following lines:
4x+16y=12
by putting the equation in slope-intercept form.
(This is what i got so far):
16y=4x+12
Slope=4

Answer by Alan3354(69443)

Find the slop and y-intercept of the following lines:
4x+16y=12
by putting the equation in slope-intercept form.
(This is what i got so far):
16y=4x+12
Slope=4
--------------
4x+16y=12
x+4y = 3
4y = -x + 3
y = (-1/4)x = 3/4
So the slope is -1/4, and the y-intercept is 3/4